### v2

# Reasoning-Aware Proxy Reward Model using Process Mining

Abstract
Recent advances in sparse reward policy gradient methods have enabled effective reinforcement learning (RL) for language models post-training. However, for reasoning tasks such as mathematical problem solving, binarized outcome rewards provide limited feedback on intermediate reasoning steps. While some studies have attempted to address this issue by estimating overall reasoning quality, it remains unclear whether these rewards are reliable proxies for the quality of stepwise reasoning. In this study, we consider reasoning as a structured process and propose TACReward, the reward model that can be seamlessly integrated into sparse reward policy gradient methods without additional human annotation costs or architectural modifications. TACReward aggregates stepwise structural deviations between teacher and policy reasoning using process mining techniques, producing a scalar output reward range of $[0,1]$ to indicate reasoning quality. Experiments on multiple mathematical reasoning benchmarks demonstrate that integrating the TACReward into sparse reward frameworks encourages the policy model to improve the structural quality of reasoning. Consequently, this leads to consistent performance improvements over existing sparse reward frameworks. Our code and model are publicly available at GitHub and HuggingFace
Large Reasoning Model, Reward, Policy Optimization, Process Mining
1 Introduction
Recent advances in post-training methods based on RL (Ouyang et al., 2022) have established a new paradigm for enhancing the alignment and reasoning capabilities of Large Reasoning Models (LRMs) (Zhang et al., 2025). Traditional policy gradient methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and its variants, have achieved significant success. To improve efficiency, Reinforced Leave-One-Out (RLOO) (Ahmadian et al., 2024) introduced a leave-one-out baseline to reduce the variance without a dedicated critic. GRPO (Shao et al., 2024) and GSPO (Zheng et al., 2025) have shifted this paradigm by adopting sparse reward signals and removing the critic network and Generalized Advantage Estimation (GAE).
Despite these advances, policy gradient methods that rely on sparse reward signals remain challenging to apply to reasoning tasks, such as mathematical problem solving. This difficulty arises from their dependence on verifiable outcome rewards (i.e., reward sparsity)  (Cui et al., 2025). This reliance on the binary and rule-based signals often produces uniform reward signals when responses within a problem group exhibit similar correctness. Consequently, this yields minimal differentiation and weakened learning gradients (Zhang and Zuo, 2025). One potential solution involves providing step-level labels such as Process Reward Models (PRMs) (Lightman et al., 2023). However, adopting such dense reward within a sparse reward framework is challenging without architectural modifications, potentially diluting the advantage of stability (Sullivan, 2025). Furthermore, defining step-level labels is inherently challenging, and human annotation remains costly (Jaech et al., 2024; Lightman et al., 2023).
Recent studies have proposed sparse reward frameworks that consider reasoning quality without annotation (Cui et al., 2025; Fan et al., 2025; Yang et al., 2025; Zhan et al., 2025), most of which are based on GRPO. However, these methods often rely on an indirect estimation or approximation of the overall reasoning process quality. This makes it challenging to consider the resulting reward signal as indicator of the quality of the individual reasoning steps. Consequently, concerns remain regarding the reliability of resulting reward signals, as it remains ambiguous whether improvements in cumulative rewards genuinely reflect enhanced logical maturity or merely the over-optimization of the surrogate metrics. Therefore, while ensuring that step-level logical integrity is crucial for authentic reasoning, an inherent dilemma arises from its sparse reward nature.
To address these challenges, this study proposes a Trace, Alignment, and Check Reward (TACReward), a novel reasoning-aware reward model that can be seamlessly integrated into sparse reward policy gradient methods for LRM post-training. TACReward treats the reasoning of LRMs as a structured process rather than a monolithic output. The core idea is that TACReward encourages the policy model to improve its degree of alignment with the teacher model reasoning process, which is considered more mature than the policy model. The degree of alignment was computed as the sparse final reward signal in the range 0 and 1 by aggregating stepwise structural deviations between the teacher and policy models. Therefore, TACReward provides proxy signals for the quality of the reasoning process from a structural alignment perspective. To this end, process mining (Van Der Aalst, 2016), a set of techniques specialized in analyzing data recording the execution of processes (i.e., event logs), is adapted and extended. Notably, these procedures require no additional human annotation.
Our proposed TACReward was evaluated using multiple mathematical benchmarks (Hendrycks et al., 2021; He et al., 2024; MAA, n.d.; Lewkowycz et al., 2022; KICE, 2025; Liu et al., 2025b). The experimental results showed that integrating TACReward into sparse reward policy gradient methods can enhance performance over RLOO, GRPO, and GSPO across all benchmarks. Notably, GSPO + TACReward achieved an average relative accuracy improvement of 89.2 %.
2 Preliminaries
2.1 Event Log & Process Mining
Figure 1: Example of mathematical reasoning event log with a single case id and the corresponding process model (i.e., petri-net) inferred from it. The boxes in the map represent activities, and the arrows indicate the process flow. The circles denote the places, which are used to model the state of the process.
Figure 2: Overview of the proposed TACReward. Given a query $x$, responses $\left\{y_{i}\right\}_{i=1}^{G}$ are sampled from the policy model $\pi_{\theta}$ and converts them into policy traces, from which $G$ process models $\left\{\mathcal{M}_{i}\right\}_{i=1}^{G}$ are discovered. The teacher’s reasoning is converted into a reference trace, and each process model $\mathcal{M}_{i}$ is evaluated against it to obtain the scalar rewards $\left\{r^{TAC}_{i}\right\}_{i=1}^{G}$. In cases where multiple responses are not generated for the same query, $G=1$.
An event log (Van Der Aalst et al., 2011) is hierarchically structured data that records the execution of a process. The formal definitions of events, traces, and logs are as follows:
Definition 2.1 (Event, Trace, and Log).
E is the event universe. An event is a tuple $e=\left(c,a,t,d_{1},\dots,d_{m}\right)\in E$, where $c$ is the case id; $a$ is the activity; $t$ is the timestamp, and $d_{1},\dots,d_{m}$ are the additional attributes. A trace $\sigma=\left\langle e_{1},e_{2},\dots,e_{n}\right\rangle\in E^{*}$ is a finite, non-empty sequence of events such that a trace contains all and only the events in the same case id. An event log $\mathbb{L}$ is a multiset of traces.
Process mining (Van Der Aalst, 2016) is a set of techniques that uses event logs to discover process models (Leemans et al., 2013), check the conformance between event log and process model (Carmona et al., 2022), and enhance or extend the models. Figure 1 presents an example of an event log and the process model derived from it. Various discovery algorithms have been proposed ranging, from foundational Alpha Miner (Van Der Aalst et al., 2011) to more robust approaches, such as inductive miners (Leemans et al., 2013).
2.2 Notations
In this study, an autoregressive LRM as a stochastic policy over sequences parameterized by $\theta$ is denoted by policy $\pi_{\theta}$. The teacher model is denoted by $\pi_{\phi}$, where $\phi$ is the set of parameters. Given an input query $x\sim\mathcal{D}$ sampled from the dataset $\mathcal{D}$, the model generates a response $y=\left\langle y_{1},y_{2},\dots,y_{\left\lvert y\right\rvert}\right\rangle$.
2.3 Sparse Reward Policy Gradient Methods for LRMs
The policy gradient methods (Sutton et al., 1999) with sparse rewards maximizes expected reward for the responses generated by the policy LRM $\pi_{\theta}$:
$\mathcal{J}(\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[R(x,y)\right]$ (1)
where $R(x,y)$ is the reward function that provides a scalar reward signal based on the quality of the response $y$ to query $x$. Using the score function estimator (Williams, 1992), the policy gradient can be expressed as
$\nabla_{\theta}\mathcal{J}(\theta)=\mathbb{E}\Big[\nabla_{\theta}\log\pi_{\theta}(y|x)\,\big(R(x,y)-b(x)\big)\Big]$ (2)
where $b(x)$ is a baseline independent of $y$, introduced to reduce the variance of the estimator. Such a baseline acts as a control variate, subtracting any function independent of $y$, and preserving the unbiasedness of the gradient estimator. For autoregressive language models, the log-likelihood factorizes over the tokens as follows:
$\log\pi_{\theta}(y|x)=\sum_{t=1}^{|y|}\log\pi_{\theta}\big(y_{t}\mid x,y_{<t}\big)$ (3)
It is often convenient to define the (Monte Carlo) advantage estimator (Sutton et al., 1998) as: $\hat{A}(x,y)\triangleq R(x,y)-b(x)$, such that Equation 2 can be expressed as $\hat{A}(x,y)$.
2.4 Group-based Policy Gradient Methods for LRMs
Recent sparse-reward post-training methods often sample multiple responses for the same query $x$ (Shao et al., 2024; Zheng et al., 2025). For each $x$, we draw a group of $G$ responses
$\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid x)$ (4)
and compute their corresponding outcome rewards $\{r_{i}\}_{i=1}^{G}\triangleq\{R\big(x,y_{i}\big)\}_{i=1}^{G}$, where $r_{i}$ is typically sparse (e.g., binary correctness) in reasoning tasks. A common baseline choice in this setting is:
$\hat{A}_{i}=r_{i}-\operatorname{mean}\left(\left\{r_{i}\right\}_{i=1}^{G}\right)$ (5)
3 TAC: Trace, Align, and Check
The proposed TACReward treats the reasoning of an LRM as a structured process rather than a monolithic output. By leveraging and Extending Techniques From process mining, TACReward evaluates the logical integrity of a reasoning trace without requiring step-level human annotations or architectural changes to the policy optimization framework. TACReward encourages a policy model to improve the degree of alignment with the teacher model’s reasoning process, which is logically more mature than the policy model. DeepSeek R1 (Guo et al., 2025) was used as the teacher model because of its strong reasoning capabilities. The core concept of TACReward is illustrated in Figure 2.
In process mining literature, conformance checking aims to evaluate how a process model can explain an entire event log that contains multiple traces (Van Der Aalst, 2016). In the proposed TACReward, this paradigm was extended to operate on individual traces. TACReward can be categorized into three main steps: 1) Trace: Formalizing the reasoning trace of both the policy and teacher models. 2) Align: Examines the alignment between the policy and reference traces. 3) Check: Evaluate how well the reasoning process of the policy model conforms to that of the teacher model.
3.1 Trace: Formalize Reasoning
Table 1: Taxonomy of 20 mathematical reasoning activities
# Activity # Activity
1 Start Problem 11 Justify Step
2 Recall Definition 12 Explore Edge Cases
3 Identify Known Results 13 Identify Contradiction
4 Formulate Strategy 14 Interpret Result
5 Apply Known Formula 15 Check Validity
6 Simplify Expression 16 Verify With Example
7 Change of Variable 17 Refine or Change Strategy
8 Evaluate Limit or Integral 18 Conclude Final Result
9 Perform Comparison 19 Recheck Original Question
10 Apply Theorem 20 End Problem
The computation of the TACReward begins by prompting the policy model to generate a reasoning process for a given problem $x$. For each problem, the model produces $G$ candidate responses, $\{y_{i}\}_{i=1}^{G}$, each containing a sequence of reasoning steps. The steps are enclosed within <think><think> tags (see Section B.1 for the generation prompt). Subsequently, these responses were formalized into reasoning traces by segmenting them into discrete steps and formatting them to conform to an event- log schema. Each fundamental units of reasoning represents a distinct activity. The taxonomy of the 20 activities is adopted based on related studies and mathematical problem-decomposition frameworks (Qin et al., 2025; Berti et al., 2025; Polya, 2014; Ritter et al., 2019). A complete list of the activities is provided in Table 1.
To extract the traces, we employed DeepSeek-V3.2 (Liu et al., 2025a), which is a powerful general-purpose language model. Let $\pi_{\psi}$ denote a general purpose model: The policy traces are formalized as:
$\left\{\sigma_{i}\right\}_{i=1}^{G}=\pi_{\psi}\left(\pi_{\theta}(\cdot\mid x)\right)$ (6)
where $\sigma_{i}$ is the trace corresponding to the $i$-th response $y_{i}$. Similarly, the reference trace $\pi_{\phi}$ is formalized as:
$\sigma^{ref}=\pi_{\psi}\left(\pi_{\phi}(\cdot\mid x)\right)$ (7)
where the teacher model generates a single response for each problem. The prompt used to formalize the trace is provided in Section B.2.
3.2 Align: Examine the Alignment
In this step, each policy trace is transformed into a process model using the Inductive Miner (IM) algorithm (Leemans et al., 2013). Given the policy traces $\{\sigma_{i}\}_{i=1}^{G}$ and reference trace $\sigma^{ref}$, we discover a process model for each policy trace by treating $\{\sigma_{i}\}$ as single-trace event log:
$\mathcal{M}_{i}=\operatorname{IM}\left(\left\{\sigma_{i}\right\}\right),\quad i=1,\ldots,G$ (8)
where $\mathcal{M}_{i}$ denotes the discovered process model (e.g., a Petri net) corresponding to the $i$th policy trace.
To establish an alignment (Van der Aalst et al., 2012) between a policy process model $\mathcal{M}_{i}$ and the reference trace $\sigma^{ref}$, we relate the moves in the model to the moves in the trace. Let $\Sigma$ be the set of activity labels and let $\gg$ denote a ”no-move” (i.e., misalignment) symbols. Given model $\mathcal{M}_{i}$, let $\beta(\mathcal{M}_{i})\subseteq\Sigma^{*}$ be the set of all possible complete executions sequences in $\mathcal{M}_{i}$.
Definition 3.1 (Alignment).
The alignment of $\sigma^{ref}$ and $\mathcal{M}_{i}$ is a sequence: $\gamma=\langle(a^{ref}_{1},a^{m}_{1}),\dots,(a^{ref}_{K},a^{m}_{K})\rangle$ where $(a^{ref}_{k},a^{m}_{k})\in(\Sigma\cup\{\gg\})\times(\Sigma\cup\{\gg\})\setminus,\{(\gg,\gg)\}$ such that the top row yields $\sigma^{ref}$ and the bottom row yields $\sigma^{\beta}\in\beta(\mathcal{M}_{i})$. The set of all these alignments is denoted by $\Gamma(\sigma^{ref},\mathcal{M}_{i})$.
To illustrate this, consider the reference trace $\sigma^{ref}=\langle A,D,B,E,G\rangle$ and the process model in Figure 1 as the policy process model $\mathcal{M}_{i}$. Two example alignments are as follows:
[…]
where $\gg$ indicates a misalignment (a log-only move or model-only move), where each column represents a vertical movement. For example, In $\gamma_{1}$, the first move $(A,A)$ is a synchronous move. The the second move $(D,\gg)$ indicates that the reference trace executes $D$ whereas the model does not. In $\gamma_{2}$, the second move $(\gg,C)$ indicates that: that the model executes $C$ that is absent from the reference trace.
To quantify the discrepancy between the reference traces $\sigma^{ref}$ and the behavior allowed by the policy process model $\mathcal{M}_{i}$, a finite cost is assigned to each move in an alignment.
Definition 3.2 (Cost Function).
Let $(a_{ref},a_{m})$ be a move in alignment $\gamma\in\Gamma(\sigma^{ref},\mathcal{M}_{i})$. The move cost function $\delta_{m}$ is defined as
$\delta_{m}(a_{ref},a_{m})=\begin{cases}0,&a_{ref}=a_{m}\neq\gg,\\ w_{L}(a_{ref}),&a_{m}=\gg,\\ w_{M}(a_{m}),&a_{ref}=\gg\end{cases}$ (9)
where $w_{L}(\cdot)\geq 0$ and $w_{M}(\cdot)\geq 0$ are non-negative weights. We use $w_{L}(\cdot)=w_{M}(\cdot)=1$ unless stated otherwise. These three cases correspond to the following: Synchronous move, move on log, and move on model, respectively.
Subsequently, the total cost of the alignment, $\delta(\gamma)$, is computed by summing the move costs over all the moves in the alignment of $i$-th model:
$\gamma_{i}^{*}=\underset{\gamma\in\Gamma(\sigma^{ref},\mathcal{M}_{i})}{\mathrm{argmin}}\,\delta(\gamma)$ (10)
3.3 Check: Evaluate the Conformance
Through the alignment step, an optimal alignment $\gamma_{i}^{*}$ is established between each policy process model $\mathcal{M}_{i}$ and the reference trace $\sigma^{ref}$. Based on this alignment, we quantify the conformance score $r^{TACReward}_{i}$, which serves as the TACReward output for each policy response. We computed the conformance by using the F1-score of fitness and precision (Van der Aalst et al., 2012).
Fitness. Fitness measures how well the process model explains the observed behavior in the the reference traces (Van Der Aalst, 2016). To obtain stable normalization, the finite the worst-case deviation cost is defined as:
$\delta_{worst}(\sigma^{ref},\mathcal{M}_{i})$ (11)
$=\sum_{e\in\sigma^{ref}}w_{L}(e)$ $\quad+\min_{\sigma^{\beta}\in\beta(\mathcal{M}_{i})}\sum_{e^{\prime}\in\sigma^{\beta}}w_{M}(e^{\prime})$
the cost of $\sigma^{ref}$ with only log moves, plus the minimum cost to complete the model using only the model moves. Subsequently, the fitness is computed as
$s_{fit}(\sigma^{ref},\mathcal{M}_{i})=1-\frac{\delta(\gamma_{i}^{*})}{\delta_{worst}(\sigma^{ref},\mathcal{M}_{i})}$ (12)
Precision. The precision assesses how often actions that are not present in the trace occur in a process model (Van Der Aalst, 2016). Let $en_{M}(e)$ be the set of enabled activities in model $\mathcal{M}_{i}$ in the state immediately before event $e$ occurs, and let $en_{L}(e)$ be the set of activities observed in $\sigma^{ref}$ in the same context (i.e., the same prefix immediately before $e$). The precision is computed as:
$s_{prec}(\sigma^{ref},\mathcal{M}_{i})=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\frac{|en_{L}(e)|}{|en_{M}(e)|},$ (13)
where $\mathcal{E}$ denotes the collection of events in the reference trace. A precision value close to 1 indicates that the model does not allow significantly more behavior than that observed in the reference trace.
Output of TACReward. Based on the computed fitness and precision, the output reward $r^{TAC}_{i}$ is defined as
$r^{TAC}_{i}=\frac{2\cdot s_{fit}(\sigma^{ref},\mathcal{M}_{i})\cdot s_{prec}(\sigma^{ref},\mathcal{M}_{i})}{s_{fit}(\sigma^{ref},\mathcal{M}_{i})+s_{prec}(\sigma^{ref},\mathcal{M}_{i})}$ (14)
where $r^{TAC}_{i}\in[0,1]$ denotes the degree of conformance between the reasoning process of the policy model and in the teacher model for the $i$th response. TACReward can be seamlessly integrated into sparse-reward policy gradient methods by directly adding its output as the original task reward. Detailed integration procedures are provided in Appendix C.
3.4 TAC as a Proxy for Step-level Reward
While $r^{TAC}_{i}$ is delivered as a single scalar value per response, $r^{TAC}_{i}$ is computed from a fine-grained step-level comparison of reasoning processes via the alignment $\gamma_{i}^{*}$. Consequently, TACReward provides a proxy for step-level supervision without requiring explicit dense annotations.
The alignment $\gamma_{i}^{*}$ decomposes the discrepancy between the policy and teacher reasoning traces into individual moves, each incurring local cost $\delta_{m}(\cdot)$. The total deviation cost performs implicit credit assignment at the reasoning transition level, even though the final reward is scalar. A log-only move $(a_{ref},\gg)$ indicates that the policy model omits the reasoning activity that is present in the teacher’s trace and corresponds to a missing logical step. Conversely, the model-only move $(\gg,a_{m})$ reflects the introduction of redundant or irrelevant reasoning activities that are not supported by the teacher’s process.
Under this decomposition, the fitness penalizes missing essential steps, whereas the precision penalizes excessive or unjustified behavior. Their harmonic mean captures both completeness and selectivity and prevents premature conclusions and unsupported over generations.
4 Experiments
Table 2: Optimization behavior of sparse-reward RL methods with TACReward (240 training steps)
MATH 500 MINERVA Olympiad AIME 2024 AIME 2025 KSAT 2025 LiveMath Bench Avg.*
Qwen2.5-7B-Instruct
PPO 43.2 25.4 35.6 10.0 6.7 36.7 7.0 22.3
[…]
RLOO 57.4 25.4 37.6 10.0 0.0 30.0 9.0 20.4
[…]
GRPO 57.4 26.1 33.3 13.3 3.3 30.0 7.0 19.9
[…]
GSPO 61.4 19.9 33.3 10.0 10.0 16.7 6.0 17.2
[…]
Table 3: Comparison with state-of-the-art RL methods using GSPO + TACReward (3000 training steps)
MATH 500 MINERVA Olympiad AIME 2024 AIME 2025 KSAT 2025 LiveMath Bench
DeepSeek-R1-Distill-Qwen-7B
BASELINE (Shao et al., 2024) 80.6 40.4 53.8 43.3 30.0 66.7 13
[…]
DeepMath (He et al., 2025b) 83.4 42.6 48.7 34.2 30.0 63.3 17
SkyWork (He et al., 2025a) 87.0 43.4 55.7 46.7 36.7 66.7 12
GRPO-LEAD (Zhang and Zuo, 2025) 84.6 47.4 52.3 40.0 26.7 76.7 13
DRGRPO (Liu et al., 2025c) 80.2 10.3 41.0 40.0 6.7 66.7 21
ExGRPO (Zhan et al., 2025) 82.8 41.0 47.9 23.3 16.7 63.3 13
Eurus-2 (Cui et al., 2025) 79.2 38.6 42.1 26.7 16.7 20.0 2
VeriThinker (Chen et al., 2025b) 80.2 34.9 46.4 30.0 13.3 76.7 14
DeepSeek-R1-Distill-Qwen-1.5B
BASELINE (Shao et al., 2024) 69.4 26.5 40.0 23.3 23.3 60.0 11
[…]
DRA-GRPO (Chen et al., 2025a) 71.0 26.1 40.4 20.0 23.3 43.3 5
Open-RS3 (Dang and Ngo, 2025) 69.8 26.8 39.7 20.0 20.0 50.0 6
STILL-3 (Min et al., 2024) 72.4 23.2 42.8 20.0 16.7 53.3 3
ExGRPO (Zhan et al., 2025) 71.2 30.9 34.4 10.0 10.0 40.0 10
Figure 3: Reward changes during GSPO + TAC training, smoothed with a 15-step moving average. Left: 7B model. Right: 1.5B model.
Figure 4: Changes in Entropy, Mean Length, TACReward for Fitness vs. Precision vs. Both (240 training steps)
4.1 Experiment Setup
Models & Baselines
We consider two categories of base models. For non-reasoning base models, we used Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct (Qwen et al., 2025), which did not were fine-tuned with RL for reasoning capabilities. For the reasoning-enhanced base models, we use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), which already possess reasoning capabilities based on distillation.
Benchmarks & Evaluation Metric
We train on the DeepMath-103k (He et al., 2025b) dataset and evaluated seven challenging mathematical reasoning benchmarks: MATH-500 (Hendrycks et al., 2021) MINERVA (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), LiveMathBench (Liu et al., 2025b) and Korean CSAT Math Calculus (KICE, 2025), and AIME 2024–2025 (MAA, n.d.). However, we did not address this problem MATH500 and AIME 2024 as primary evidence because of potential contamination (Wu et al., 2025). Therefore, the results for MATH-500 and AIME 2024 are reported for reference only, and the system prompts that was previously mentioned, was not used. Instead, all of the main quantitative comparisons and conclusions are obtained from the remaining benchmarks. All evaluations used the Pass@1 metric temperature = 0.6, top_p = 0.95, and max_tokens = 16384, following the flashinfer (Ye et al., 2025) framework. The answers are extracted using the DeepMath Evaluation Library  (He et al., 2025b).
RL Settings
All experiments are conducted using the TRL framework (von Werra et al., 2020) on $4\times$ NVIDIA H200 GPUs. We employed the DeepSeek API (Guo et al., 2025) as the reward model that provides both accuracy and conformance reward signals. We used the AdamW optimizer for optimization at a constant learning rate of $1\times 10^{-6}$. The training was performed with a global batch size of 128, utilizing micro-batch sizes of 8 for the 1.5B models and 4 for 7B models, respectively. We set the KL divergence coefficient to zero for unconstrained policy updates. The details of hyperparameters are provided in Appendix A.
4.2 Foundational Policy Gradients Methods with TACReward
Table 2 compares four representative sparse-reward policy optimization methods—PPO, RLOO, GRPO, and GSPO, with and without TACReward-under a fixed training budget of 240 optimization steps on Qwen2.5-7B-Instruct. This setting characterizes early stage optimization rather than converged performance. To better assess the performance, we compute the average score (Avg. *) by excluding MATH500 and AIME 2024 and instead focus on the for the remaining five benchmarks.
Strong Synergy with GSPO.
The most significant improvement was observed when TACReward was combined with GSPO. Although GSPO alone underperforms in this early stage regime, GSPO + TAC increases Avg. * from 17.2 to 32.5 (+89.2%). We attribute this synergy to a granularity match between the optimization objective and reward signal. GSPO defines the importance ratio and updates it at the sequence level, effectively treating each sampled response as a single optimization unit. TACReward aggregates the step-level deviations revealed by the trace alignment into a single sequence-level conformance score that represents the reasoning quality of one response. Conseqently, both components operate under a coherent unit of credit assignment—one reward per response. This allows TACReward to provide more informative and a well-aligned training signal for GSPO and can be exploited more effectively than in token-level or step-mismatched objectives.
General Improvements with RLOO and GRPO.
TACReward provides consistent improvements for RLOO and GRPO, with relative gains of +6.1% and +12.7%, respectively, in Avg. This suggests a broad compatibility with variance-reduced and group-based objectives. Stronger baselines are less sensitive to the additional variance introduced by proxy rewards.
Limited gains with PPO.
Adding TACReward to PPO yielded limited gains and slightly reduced the overall average score (Avg. *) despite improvements in some benchmarks. This behavior can be explained by the inherent characteristics of PPO. These characteristics aim to dense reward settings, and can be adapted to sparse rewards. In this study, PPO was trained under sparse reward regime, in which sensitivity to reward variance and credit assignment becomes more pronounced. Appendix E provides visualizations of reasoning processes with and without TACReward during GSPO training.
4.3 Comparison with State-of-the-Art Models
Based on the controlled comparison in Table 2, we selected GSPO + TACReward as the most effective configuration and optimized it further for evaluation. Subsequently, we compared this configuration with the recent state-of-the-art RL methods. Table 3 compares GSPO + TAC with the recent state-of-the-art RL methods for two distilled backbone models of different sizes. Training was conducted using 3000 training steps. For both model sizes, GSPO + TAC exhibited strong or state-of-the-art performances across the benchmarks. With the 7B backbone, it outperformed all datasets, particularly Olympiad and KSAT 2025, and remained competitive (often better) at 1.5B.
Reward Dynamics during Training.
Figure 3 shows that TACReward steadily increases during training, indicating a progressive better conformance between the policy’s reasoning process and teacher’s traces. For both the model sizes, the model learns to produce more accurate answers while more closely aligning the reasoning process with the teacher. Appendix F provides visualizations of reasoning processes as training progresses, indicating that the model’s reasoning steps became more structured and aligned with those of the teacher over time.
Proof of Concept for Using Both Fitness and Precision.
Based on the definitions in (Van Der Aalst, 2016), fitness penalizes the missing essential steps, whereas precision penalizes excessive or unjustified behavior. Through Figure 4, we find that this definition aligns exactly with the expected behavior patterns that we want to encourage during training.
Table 4: Ablation Study on Reward Components: Fitness vs. Precision vs. Both (240 training steps)
Minerva Olympiad AIME 2025 KSAT 2025 LiveMath Bench
w/ Fit 30.9 30.5 6.7 43.3 10
w/ Prec 35.7 35.7 3.3 46.7 12
w/ Both 37.1 38.2 10.0 53.3 15
First, when only fitness was used, we expected the model to generate as many reasoning steps as possible and to employ a wide variety of activities to avoid penalization for missing essential steps. This is confirmed in Figure 4, where using only fitness (green) results in longer reasoning chains (mean length). However, when only precision is used (purple), the model becomes conservative, generating fewer reasoning steps and using fewer activities to minimize excessive behavior. This is indicated by the shorter mean length and higher entropy. Both components (red) balance these two tendencies, yielding the highest confidence (lowest entropy) and moderate length. Appendix G provides visualizations of reasoning processes, where the process models extracted from the same problem visually demonstrate this assumption. Table 4 further confirms that using either component alone leads to a degraded performance across benchmarks.
5 Discussion
5.1 Design Rationale and Methodological Choices
Role of Mathematical Taxonomy.
The Mathematical Reasoning Taxonomy (MRT) provides a fixed set of reasoning activities for formalizing raw reasoning responses into comparable traces.
Table 5: Ablation Study on the Effect of Mathematical Reasoning Taxonomy (MRT) (240 training steps)
Minerva Olympiad AIME 2025 KSAT 2025 LiveMath Bench
w/o MRT 37.1 35.6 3.3 43.3 9
w/ MRT 37.1 38.2 10.0 53.3 15
Table 5 indicates that removing MRT degrades performance, with larger drops on multistep benchmarks (e.g., AIME 2025 and KSAT 2025). In this setting, the rewards become less sensitive to structural differences and are dominated by surface variation without an explicit taxonomy. Appendix D provides further discussions on the design rationale behind TACReward.
5.2 Limitations
Computational Overhead.
TACReward includes additional computations for trace formalization, process discovery, and conformance checking.
Table 6: Per-Step Training Time (s) Across Algorithms With/Without TACReward
PPO RLOO GRPO GSPO
w/o TAC 8.0 8.5 8.1 10.9
w/ TAC 39.7 37.9 64.3 67.2
As shown in Table 6, TACReward increases the per-step training time across all algorithms. This overhead is the main trade-off in the proposed approach.
6 Related Works
Recent studies have explored approaches that consider the quality of reasoning processes in a sparse reward policy gradient methods for LRMs (Cui et al., 2025; Fan et al., 2025; Yang et al., 2025; Zhan et al., 2025). However, these methods rely on an indirect estimation or approximation of the overall reasoning quality. For example, metrics such as the final answer correctness, model confidence (e.g., entropy) (Zhan et al., 2025), preferences for optimized-degraded reasoning pairs (Fan et al., 2025), expectations of tree-sampled continuations (Yang et al., 2025), and Process rewards derived from outcome supervision(Cui et al., 2025) were used to estimate the overall reasoning quality. Although these methods capture reasoning quality, it remains challenging to consider the resulting reward signal as reflecting the quality of the individual reasoning steps. In contrast, TACReward is computed based on the step-level deviations in the reasoning process. Consequently, it serves as a proxy for step-level reasoning quality.
7 Conclusion
Sparse reward policy gradient methods are challenging to incorporate into reasoning tasks, particularly when learning which reasoning steps to perform and in which order. Moreover, providing such annotations is costly, and their quality can significantly affect training outcomes. In this study, we proposed TACReward, a reasoning-aware proxy reward that can be seamlessly integrated into sparse reward policy gradient methods for LRM post-training without additional annotations or architectural changes. Instead of teaching the model’s exact intermediate actions, TACReward measures the structural alignment between the reasoning processes of a policy model and that of a more logically mature teacher. The results showed that accounting for structural similarity improves the sparse reward performance and optimization behaviors induced by fitness and precision match their intended effects. Therefore, TACReward effectively captures key structural properties of reasoning. However, TACReward introduces additional computational overhead. Future studies must focus on reducing API dependence and improving the efficiency.
Impact Statement
This study proposes a TACReward for advancing the field of language model post-training. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Details of Hyperparameters
We trained the policy model using the TRL framework with DeepSpeed ZeRO-3 optimization. Table 7 summarizes the key hyperparameters and distributed training configuration used in experiments 4. All other parameters followed the default settings in TRL GRPOTrainer.
Table 7: Hyperparameters and Training Configuration for TAC Reward with GSPO
Module Parameter Value Description
Data dataset_name DeepMath-103k Training dataset.
max_prompt_length 1024 Maximum input prompt length.
max_completion_length 16384 Maximum response length.
per_device_train_batch_size 2 Batch size per device.
Model torch_dtype bfloat16 Model precision.
attn_implementation flash_attention_2 Attention implementation.
use_vllm True Enable vLLM for inference.
vllm_mode colocate vLLM execution mode.
Optimizer learning_rate $1\times 10^{-6}$ Learning rate for policy optimizer.
max_steps 1000 Total number of training steps.
gradient_accumulation_steps 4 Gradient accumulation steps.
optimizer AdamW Optimizer type (TRL default).
GRPO num_generations 8 Number of rollouts per prompt.
importance_sampling_level sequence Importance sampling granularity.
beta 0.0 KL penalty coefficient.
epsilon 0.2 Clipping parameter (default).
DeepSpeed distributed_type DEEPSPEED Distributed training backend.
zero_stage 3 ZeRO optimization stage.
zero3_save_16bit_model True Save model in 16-bit precision.
mixed_precision bf16 Mixed precision training.
num_processes 4 Number of GPU processes.
Appendix B Prompts for Model Training
B.1 Prompt for Generation
[…]
B.2 Prompt for Formalizing Trace
Prompt Template for Trace Formalization System Role: You are an Expert Mathematical Problem Solver and Process Mining Assistant. Your goal is to convert unstructured logical reasoning steps into a structured, generalized CSV event log ready for process mining analysis. Guidelines for Generalization:
You must select the Activity label that best describes the intent of the reasoning step. Use the definitions below to ensure consistency:
1. Start Problem 2. Recall Definition 3. Identify Known Results 4. Formulate Strategy 5. Apply Known Formula 6. Simplify Expression 7. Change of Variable 8. Evaluate Limit or Integral 9. Perform Comparison 10. Apply Theorem 11. Justify Step 12. Explore Edge Cases 13. Identify Contradiction 14. Interpret Result 15. Check Validity 16. Verify With Example 17. Refine or Change Strategy 18. Conclude Final Result 19. Recheck Original Question 20. End Problem Absolute Constraints: NEVER invent new activities. Use ONLY the 20 labels listed above. ALWAYS select exactly one of the above labels for the Activity column. Use each activity label according to its abstract logical purpose. Each step should reflect one unit of reasoning with a concise description. Process the input line-by-line: output exactly one CSV row per non-empty line. Example Output Format: Case ID Step Activity Description 1 01 Start Problem Problem: Find function g(x) such that... 1 02 Formulate Strategy Break down the problem into constraints 1 03 Recall Definition Remember that $x^{n}e^{-x}$ is bounded... 1 04 Perform Comparison Compare integrals... 1 05 Apply Known Formula Use integral of $e^{-x}$ from 0 to 1 1 06 Simplify Expression Solve for c such that total = 1 1 07 Check Validity Ensure inequality is strict 1 08 Conclude Final Result Final answer: $g(x)=e^{-x}+1/e$ 1 09 End Problem Problem solved with valid solution Input:
{problem}
Return your output only in CSV format starting with the header.
Appendix C Interating TACReward with Sparse Reward Policy Gradients Methods
In this section, we demonstrate how TACReward can be easily integrated into the foundational sparse reward policy gradient methods: PPO, RLOO, GRPO, and GSPO. In this section, we show how TACReward can be easily integrated into foundational sparse reward policy gradient methods: PPO, RLOO, GRPO, and GSPO.
Notations
TACReward is denoted as $\left\{r^{TAC}(x,y_{i})\right\}_{i=1}^{G}$. We set $\left\{r^{acc}(x,y_{i})\right\}_{i=1}^{G}$ as the accuracy reward and $\left\{r^{fmt}(x,y_{i})\right\}_{i=1}^{G}$ as follows: thinking format reward for each response $\left\{y_{i}\right\}_{i=1}^{G}$. The total reward for each query-response pair is denoted as $\left\{R(x,y_{i})\right\}_{i=1}^{G}$ as described in Section 2. In cases where multiple responses are not generated for the same query $G=1$ and are simply denoted as $R(x,y)$.
C.1 Proximal Policy Optimization (PPO) with TACReward
PPO (Schulman et al., 2017) generates a single response for each input query (i.e., $G=1$) and update the policy model based on the rewards obtained from the response. For an input $x\sim\mathcal{D}$ and a sampled response $y=\left\langle y_{1},y_{2},\dots,y_{\left\lvert y\right\rvert}\right\rangle\sim\pi_{\theta_{\mathrm{old}}}(\cdot|x)$, TACReward can be incorporated into PPO by simply adding it as follows:
$R(x,y)=r^{acc}(x,y)+r^{fmt}(x,y)+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r^{TAC}(x,y)}-\beta\sum_{t=1}^{\left\lvert y\right\rvert}\bigg(\log\frac{\pi_{\theta}(y_{t}|x_{t})}{\pi_{\mathrm{ref}}(y_{t}|x_{t})}\bigg),$ (15)
where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\beta>0$ controls the KL regularization strength. We define a sequence-level advantage with baseline $b(x)$ and broadcast it to all timesteps:
$A_{t}\triangleq A(x,y)=R(x,y)-b(x),\qquad t=1,\dots,\left\lvert y\right\rvert$, (16)
where $b(x)$ is a baseline (e.g., a learned value function $V_{\phi}(x)$) to reduce variance. Let $\omega_{t}(\theta)=\frac{\pi_{\theta}(y_{t}|x_{t})}{\pi_{\theta_{\mathrm{old}}}(y_{t}|x_{t})}$. Subsequently, the PPO clipped objective is:
$\mathcal{L}_{\mathrm{PPO}}(\theta)=\mathbb{E}\!\left[\sum_{t=1}^{\left\lvert y\right\rvert}\min\!\Big(\omega_{t}(\theta)A_{t},\;\mathrm{clip}(\omega_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\Big)\right].$ (17)
C.2 REINFORCE with Leave-One-Out (RLOO) with TACReward
RLOO (Ahmadian et al., 2024) generates a group of $G$ responses for each input query (i.e., $G>1$) and updates the policy using a leave-one-out baseline. Given an input $x\sim\mathcal{D}$ and set of sampled responses $\{y_{i}=\langle y_{i,1},\dots,y_{i,|y_{i}|}\rangle\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|x)$, TACReward can be integrated into RLOO by adding it to the reward for each response as follows:
$R(x,y_{i})=r^{acc}(x,y_{i})+r^{fmt}(x,y_{i})+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r^{TAC}(x,y_{i})}-\beta\sum_{t=1}^{|y_{i}|}\bigg(\log\frac{\pi_{\theta}(y_{i,t}|x_{i,t})}{\pi_{\mathrm{ref}}(y_{i,t}|x_{i,t})}\bigg),\qquad i=1,\dots,G,$ (18)
where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\beta>0$ controls the KL regularization strength.
RLOO uses a leave-one-out baseline computed from other samples in the group:
$b_{i}(x)=\frac{1}{G-1}\sum_{j\neq i}R(x,y_{j}),\qquad A_{i}\triangleq A(x,y_{i})=R(x,y_{i})-b_{i}(x).$ (19)
Using the score function estimator, the RLOO policy gradient is
$\nabla_{\theta}\mathcal{J}(\theta)=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Big(\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\Big)\,A_{i}\Bigg],\quad\text{where}\quad\log\pi_{\theta}(y_{i}|x)=\sum_{t=1}^{|y_{i}|}\log\pi_{\theta}(y_{i,t}|x_{i,t}).$ (20)
C.3 Group-based Policy Optimization with TACReward
GRPO and GSPO bypass the need for the value model by computing the group-relative advantage of each response within a response group for the same query. Partially borrowed from (Liu, 2025), the group-relative advantages for the responses $\left\{y\right\}^{G}_{i=1}$ are computed as:
$\hat{A}_{i,t}=\hat{A}_{i}=\frac{R(x,y_{i})-\operatorname{mean}\left(\left\{R(x,y_{i})_{i=1}^{G}\right\}\right)}{\operatorname{std}\left(\left\{R(x,y_{i})_{i=1}^{G}\right\}\right)}$ (21)
where all the tokens in $y_{i}$ share the same advantage as $\hat{A}_{i}$. GRPO then defines the objective as
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left\lvert y_{i}\right\rvert}\sum_{t=1}^{\left\lvert y_{i}\right\rvert}\quad\bigg[\operatorname{min}\bigg(\omega_{i,t}(\theta)\hat{A}_{i,t},\operatorname{clip}\left(\omega_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\bigg)\bigg]\bigg]$ (22)
where $\epsilon$ is a hyperparameter for clipping range, and the importance ratio $\omega_{i,t}(\theta)$ is $\frac{\pi_{\theta}\left(y_{i,t}|x,y_{i,<t}\right)}{\pi_{\theta_{old}}\left(y_{i,t}|x,y_{i,<t}\right)}$ The GSPO simplifies this objective as follows:
$\mathcal{J}_{GSPO}(\theta)=\mathbb{E}\bigg[\frac{1}{G}\sum_{i=1}^{G}\operatorname{min}\bigg(s_{i}(\theta)\hat{A}_{i},\operatorname{clip}\left(s_{i}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\bigg)\bigg]$ (23)
where the importance ratio $s_{i}(\theta)$ is $\bigg(\frac{\pi_{\theta}\left(y_{i}|x\right)}{\pi_{\theta_{old}}\left(y_{i}|x\right)}\bigg)^{\frac{1}{\left\lvert y_{i}\right\rvert}}$ In both GRPO and GSPO, TACReward can be integrated by simply adding it to the reward for each response as follows:
$R(x,y_{i})=r^{acc}(x,y_{i})+r^{fmt}(x,y_{i})+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r^{TAC}(x,y_{i})}\qquad i=1,\dots,G$ (24)
Appendix D Additional Discussion
D.1 Why Single-Trace Can Be Effective.
Although process discovery is traditionally applied to multitrace event logs, our use case differs fundamentally from classical approaches. We aimed to infer a global reasoning process and construct a local structural abstraction that captures the permissible reasoning transitions implied by a single policy trace. For each single trace, a minimal process model was constructed that could be generalized beyond the observed sequence, encoding the ordering constraints, optional branches, and loops in the reasoning trace. This abstraction enabled conformance checking, allowing us to penalize structurally invalid reasoning behaviors.
D.2 Why Response-Level Comparison Is Insufficient.
A natural question is whether trace formalization can be skipped, and whether reasoning responses can be directly compared as sequences. Given two token sequences, $y$ and $y^{\prime}$, the classical edit distance (Wagner and Fischer, 1974) is
$d_{\mathrm{ed}}(i,j)=\min\begin{cases}d_{\mathrm{ed}}(i-1,j)+1,\\ d_{\mathrm{ed}}(i,j-1)+1,\\ d_{\mathrm{ed}}(i-1,j-1)+\mathbb{I}[y_{i}\neq y^{\prime}_{j}]\end{cases}$ (25)
that measures surface-level similarity via insertions, deletions, and substitutions. However, $d_{\mathrm{ed}}$ measures only surface edit operations on raw text, and cannot reliably capture the structural properties of reasoning (e.g., missing or spurious transitions) that are distinct from benign re-orderings (Song et al., 2024). Trace formalization separates reasoning structure from linguistic realization, enabling a meaningful structural comparison.
D.3 On Imperfect Reference Traces.
In this study, our approach does not assume that the reference model is correct. This assumption is shared using several existing post training methods, including GRPO and GSPO. These methods rely on reward models or KL-divergence regularization with respect to a reference policy, despite these references being imperfect (Ouyang et al., 2022). The reference model must be more mature than the policy being optimized, providing a relatively stable and coherent reasoning process. Under this assumption, conformance-based rewards suppress early stage randomness and structurally invalid reasoning, whereas outcome-based rewards continue to guide convergence as the training progresses.
D.4 Applicability to Other Domains.
This study focused on mathematical reasoning; however, the proposed framework is not inherently domain-specific. Applying TACReward to other domains would require defining an appropriate set of reasoning activities to formalize the traces. Automating the discovery or adaptation of such taxonomic activities is a crucial future research direction.
Appendix E Reasoning Process Visualization with and without TACReward
Appendix F Reasoning Process Visualization across Training Progress
Appendix G Reasoning Process Visualization under Fitness, Precision, and F1-Score